OpenStudy (anonymous):

At a raffle, 1500 tickets are sold at $2 each for four prizes of $500, $250, $150, and $75. You buy one ticket. What is the expected value of your purchase?

OpenStudy (kropot72):

If you win, the profit is the value of the prize minus the cost of the ticket. If you lose, the loss is the cost of the ticket. \[E(value)=(-2\times\frac{1499}{1500})+\frac{498}{1500}+\frac{248}{1500}+\frac{148}{1500}+\frac{73}{1500}=you\ can\ calculate\]